# The Mittag-Leffler theorem for proper minimal surfaces and directed meromorphic curves

### v1

@@ 3.  The Mergelyan theorem for meromorphic $A$-immersions
If in Theorem 3.1 we additionally assume that $A_{*}=A\setminus\{0\}$ is an Oka manifold, $S$ is connected and the inclusion $S\hookrightarrow M$ induces an isomorphism $H_{1}(S,\mathbb{Z})\to H_{1}(M,\mathbb{Z})$ of the homology groups, then there is a full meromorphic $A$-immersion $G\in\mathscr{I}_{A}(M|E)$ satisfying (i) and (ii) in the statement of Theorem 3.1. Indeed, for this we adapt the proof of Lemma 3.3 as follows. Firstly, using Oka approximation [16, Theorem 5.4.4] (see also [6, Theorem 1.13.3]), we approximate the map $\tilde{f}_{0}\in\mathscr{A}(S,A_{*})$ in (3.2) uniformly on $S$ by $h_{0}\in\mathscr{O}(M,A_{*})$ such that $h_{0}-\tilde{f}_{0}$ vanishes to order $r_{0}+1$ on $E\cup(\Lambda\cap\mathring{S})$. Secondly, by the Mergelyan approximation theorem on admissible sets [15, Theorem 16] (see also [6, Theorem 1.12.11]), we approximate each $h_{i,j}\in\mathscr{O}(S)$ in (3.10) uniformly on $S$ by $\hat{h}_{i,j}\in\mathscr{O}(M)$ such that $\hat{h}_{i,j}-h_{i,j}$ vanishes to order $r_{0}+1$ on $E\cup(\Lambda\cap\mathring{S})$. Write $\widehat{H}=(\hat{h}_{i,j})_{i,j}$. The map $\Phi_{\widehat{H},h_{0}}\colon U\times M\to A_{*}$, defined as (3.12) but using $h_{0}$ and $\hat{h}_{i,j}$ from above, satisfies $\Phi_{\widehat{H},h_{0}}(0,\cdot)=h_{0}$, $\Phi_{\widehat{H},h_{0}}\approx\Phi_{H,\tilde{f}_{0}}$ uniformly on $U\times S$, and their difference vanishes to order $r_{0}+1$ on $U\times(E\cup(\Lambda\cap\mathring{S}))$. By the same argument as before, provided that all approximations are sufficiently close, the implicit function theorem furnishes us with a parameter $\zeta_{0}\in U$ close to the origin, such that the full map $\tilde{f}:=\Phi_{\widehat{H},h_{0}}(\zeta_{0},\cdot)/g\in\mathscr{O}_{\infty}(M|E,A_{*})$ satisfies $\mathscr{P}(\tilde{f})=\mathscr{P}(f)=0$, $\tilde{f}\approx f$ on $S$, and the difference $\tilde{f}-f$ vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$. The immersion $G$ is then obtained as in the proof of Theorem 3.1.
@@
Let $F$ be as in the statement. Set $\varepsilon_{0}:=\varepsilon/2$ and fix a nowhere vanishing holomorphic $1$-form $\theta$ on $M$. Remark 3.6 provides a neighbourhood $U\subset M$ of $S$ with $H_{1}(S,\mathbb{Z})\overset{\sim}{=}H_{1}(U,\mathbb{Z})$, and a full meromorphic $A$-immersion $F^{0}\in\mathscr{I}_{A}(U|E)$ such that $F^{0}-F$ is continuous on $S$, vanishes on $\Lambda\setminus\mathring{S}$, vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$ and $||F^{0}-F||_{S}<\varepsilon_{0}$.
Choose a strongly subharmonic Morse exhaustion function $\tau\colon M\to\mathbb{R}$ with $S\subset\{\tau<0\}\subset\{\tau\leq 0\}\subset U$; see e.g. [6, Proposition 1.12.5]. We may assume that 00 is a regular value of $\tau$ and that every level set $\{\tau=c\}$, $c>0$, contains at most one critical point of $\tau$. Choose a strictly increasing sequence $0=c_{0}<c_{1}<c_{2}<\dots$ with $\lim_{j\to\infty}c_{j}=+\infty$ and such that for every $j\in\mathbb{Z}_{+}$, $c_{j}$ is a regular value of $\tau$, and the set $A_{j}=\{c_{j-1}<\tau<c_{j}\}$ ($j\in\mathbb{N}$) contains at most one critical point of $\tau$. Set $M_{j}=\{\tau\leq c_{j}\}$, $j\in\mathbb{Z}_{+}$. We shall inductively construct a sequence of full meromorphic $A$-immersions $\{F^{j}\colon M_{j}\setminus E\to\mathbb{C}^{n}\}_{j}$, $F^{j}\in\mathscr{I}_{A}(M_{j}|E)$, and a decreasing sequence of positive numbers $\{\varepsilon_{j}\}_{j}$, satisfying the following conditions for all $j\in\mathbb{N}$.
@@
    $20<\varepsilon_{j}<\varepsilon_{j-1}/$ 2, and if $G\colon M\setminus E\to\mathbb{C}^{n}$ is holomorphic with $||G-F^{j}||_{M_{j}\setminus E}<\varepsilon_{j-1}$, then $G|_{M_{j-1}\setminus E}\colon M_{j-1}\setminus E\to\mathbb{C}^{n}$ is an immersion.
If such sequences exist, then the sequence $\{F^{j}\}_{j}$ converges uniformly on compact sets to the limit map $G:=\lim_{j\to\infty}F^{j}\colon M\setminus E\to\mathbb{C}^{n}$, which is a full meromorphic $A$-immersion of class $\mathscr{I}_{A}(M|E)$, the difference $G-F$ extends continuously to $S$, vanishes on $\Lambda\setminus\mathring{S}$ and vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$, and $||G-F||_{S}<\varepsilon$. So, $G$ satisfies the conclusion of the theorem.
@@
Case 2: The domain $A_{j}=\{c_{j-1}<\tau<c_{j}\}$ contains a critical point $p$ of $\tau$. By assumption, $p$ is the unique critical point of $\tau$ on $A_{j}$ and it is a Morse point with Morse index either 00 or $1$. If the Morse index of $\tau$ at $p$ equals $1$, the change of topology is described by attaching to $M_{j-1}$ a smoothly embedded oriented arc $\Gamma_{j}\subset A_{j}\cup bM_{j-1}$ with an initial point $p_{j}\in bM_{j-1}$ and a final point $q_{j}\in bM_{j-1}$, and such that $\Gamma_{j}$ intersects $bM_{j-1}$ transversely and is otherwise contained in $A_{j}$. The set $S_{j}:=M_{j-1}\cup\Gamma_{j}$ is admissible in $M_{j}$ and has the same topology as $M_{j}$. If the points $p_{j},q_{j}$ belong to the same connected component of $M_{j-1}$, a new nontrivial curve appears in the homology, however, if they belong to different components of $M_{j-1}$, we have that $H_{1}(M_{j-1},\mathbb{Z})\overset{\sim}{=}H_{1}(S_{j},\mathbb{Z})$. In either case, we proceed as follows. By [6, Lemma 3.5.4], we extend $f^{j-1}$ smoothly to the arc $\Gamma_{j}$ such that it still maps to $A_{*}$ and satisfies
@@
Otherwise, the Morse index of $\tau$ at $p$ equals 00. This time, a new connected component of $\{\tau\leq c\}$ appears at $p$ when $c$ passes $\tau(p)$. Let $\Delta\subset A_{j}$ denote a smoothly bounded disc around $p$. We extend $F^{j-1}$ to $S_{j}:=M_{j-1}\cup\Delta$ by defining any full holomorphic $A$-immersion on $\Delta$. Since $S_{j}$ is admissible in $M_{j}$ and has the same topology as $M_{j}$, Remark 3.6 applied to each component of $S_{j}$ furnishes a suitable map $F^{j}\in\mathscr{I}_{A}(M_{j}|E)$. This closes the induction and completes the proof. ∎
@@ 4.  Desingularizing meromorphic $A$-immersions
In this section we prove the following more precise version of Corollary 1.7 (b) for removing self-intersections of directed meromorphic immersions. We shall later use it in proofs of Theorems 5.1 and 7.1 in order to ensure the embeddedness.
@@
If $M$ is a compact bordered Riemann surface, $E\subset\mathring{M}$ is a finite subset, and $A\subset\mathbb{C}^{n}$ $(n\geq 3)$ is a closed irreducible conical subvariety which is not contained in any hyperplane and is smooth away from 00, then every meromorphic $A$-immersion $F\colon M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ can be approximated uniformly on $M\setminus E$ by meromorphic $A$-embeddings $G\in\mathscr{I}_{A}(M|E)$ such that the difference map $G-F$ is continuous on $M$ and vanishes to any given finite order on $E$.
@@
We emphasize that the approximation by embeddings is uniform in the non-compact surface $M\setminus E$. We establish the theorem by a recursive application of the following result, which gives the technical tool to overcome this difficulty.
@@
Let $M$, $E$ and $A$ be as in Theorem 4.1. Given an open neighbourhood $\Delta\subset\mathring{M}$ of $E$, every meromorphic $A$-immersion $F\colon M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ can be approximated uniformly on $M\setminus E$ by a meromorphic $A$-immersion $G\in\mathscr{I}_{A}(M|E)$ such that $G|_{M\setminus\Delta}:M\setminus\Delta\to\mathbb{C}^{n}$ is injective, the difference map $G-F$ is continuous on $M$ and vanishes to any given finite order on $E$.
@@
The key condition in the lemma is that, while the self-intersections are eliminated only from a compact subset, namely, $M\setminus\Delta$, the approximation is uniform on $M\setminus E$.
@@
(4.1) $\bigcup_{j\geq 1}(M\setminus\Delta_{j})=M\setminus E.$
@@
    $F^{j}-F^{j-1}$ is continuous on $\widetilde{M}$ and $||F^{j}-F^{j-1}||_{\widetilde{M}}<\varepsilon_{j-1}$.
@@
    $20<\varepsilon_{j}<\varepsilon_{j-1}/$ 2, and if $G\colon\widetilde{M}\setminus E\to\mathbb{C}^{n}$ is a holomorphic map with $||G-F^{j}||_{\widetilde{M}\setminus E}<2\varepsilon_{j}$, then $G$ is an embedding on $M\setminus\Delta_{j}$.
@@
By (1j), (2j), (4j) and the properties of $\Delta_{j}$, there exists the limit map $G=\lim_{j\to\infty}F^{j}\colon\widetilde{M}\setminus E\to\mathbb{C}^{n}$, which is meromorphic on $\widetilde{M}$ with effective poles precisely in $E$ and satisfies that $G-F$ is continuous on $\widetilde{M}$. In fact, for each $j\geq 0$ we have that $G-F^{j}$ is continuous on $\widetilde{M}$ and
@@
As a consequence, by property (4j), $G\in\mathscr{I}_{A}(M|E)$ and is injective on $M\setminus\Delta_{j}$ for all $j\geq 1$. By (4.1), it turns out that $G$ is a meromorphic $A$-embedding on $M$. Finally, condition (2j) ensure that $G-F$ vanishes to the given finite order on $E\cup\Lambda$, whereas the approximation of $F$ by $G$ is guaranteed by (4.2), closing the proof. ∎
@@
The proof follows closely the arguments in that of [4, Theorem 2.4] and [6, Theorem 3.4.1] for the case $E=\varnothing$. We shall only point out the modifications which are necessary to deal with the poles and explain the first part of the lemma; the second part concerning interpolation on $\Lambda$ is established exactly as in the proof of [6, Theorem 3.4.1] and we leave it out.
@@
Let $D_{\widetilde{M}}=\{(p,p)\colon p\in\widetilde{M}\}$ be the diagonal of $\widetilde{M}\times\widetilde{M}$. Since $F$ is an immersion on $M\setminus E$ and $\widetilde{M}\subset M\setminus E$ is compact, there exists an open neighbourhood $U\subset\widetilde{M}\times\widetilde{M}$ of $D_{\widetilde{M}}$ such that $\delta F$ does not take the value $0\in\mathbb{C}^{n}$ on $\overline{U}\setminus D_{\widetilde{M}}$. Our aim is to find a meromorphic $A$-immersion $G\colon M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ such that $G-F$ is continuous and arbitrarily close to 00 uniformly on $M$, $G-F$ vanishes to the given finite order on $E$, and the difference map $\delta G$ is transverse to $0\in\mathbb{C}^{n}$ on $(\widetilde{M}\times\widetilde{M})\setminus U$. By dimension reasons and the Abraham transversality argument [1], provided that $G$ is close enough to $F$ on $\overline{U}$, then $G$ results to be injective on $\widetilde{M}=M\setminus\Delta$. For this, we need to find a neighbourhood $\Omega\subset\mathbb{C}^{N}$ of $0\in\mathbb{C}^{N}$ (for some large $N\in\mathbb{N}$) and a holomorphic map $H\colon(M\setminus E)\times\Omega\to\mathbb{C}^{n}$ meeting the following.
@@
    $\partial_{t}|_{t=0}\delta H(p,q,t)\colon\mathbb{C}^{N}\to\mathbb{C}^{n}$
@@
That being the case, the map $G=H(\cdot,t)$ satisfies the conclusion for almost every parameter $t\in\Omega$ close enough to 00.
@@
For any pair of points $p,q\in M\setminus E$, $p\neq q$, there is a deformation family $H=H^{(p,q)}(\cdot,t)$ as above, $t\in\mathbb{C}^{n}$, such that $\partial_{t}|_{t=0}\delta H(p,q,t):\mathbb{C}^{n}\to\mathbb{C}^{n}$ is an isomorphism.
@@
(cf. [4, Eq. (6.6)]), it follows that $H_{F}(\cdot,0)=F$ and $H_{F}(\cdot,t):M\setminus E\to\mathbb{C}^{n}$ is a well-defined meromorphic $A$-immersion of class $\mathscr{I}_{A}(M|E)$ such that $H_{F}(\cdot,t)-F$ is continuous on $M$ and vanishes to the given finite order on $E$ for every $t\in\mathbb{C}^{n}$ sufficiently close to 00. Furthermore, we may ensure that the vectors
@@
defined analogously to those in the proofs of [4, Theorem 2.4] and [6, Theorem 3.4.1], satisfies $H(\cdot,0)=F$ and its difference map $\delta H$ (4.3) is submersive on $(\widetilde{M}\times\widetilde{M})\setminus U$ for all $t=(t^{1},\dots,t^{m})$ close to $0\in\mathbb{C}^{N}$ (here, $N=mn$). Moreover, $H(\cdot,t)$ is a meromorphic $A$-immersion of class $\mathscr{I}_{A}(M|E)$, $H(\cdot,t)-F$ is continuous on $M$ and vanishes to any given finite order on $E$ for every $t\in\Omega\subset\mathbb{C}^{N}$ (where $\Omega\subset\mathbb{C}^{N}$ is a neighbourhood of 00 which depends on all $\Omega_{i}$).

### v2

@@ 3. The Mergelyan theorem for meromorphic $A$-immersions
If in Theorem 3.1 we additionally assume that $A_{*}=A\setminus\{0\}$ is an Oka manifold, $S$ is connected and the inclusion $S\hookrightarrow M$ induces an isomorphism $H_{1}(S,\mathbb{Z})\to H_{1}(M,\mathbb{Z})$ of the homology groups, then there is a full meromorphic $A$-immersion $G\in\mathscr{I}_{A}(M|E)$ satisfying (i) and (ii) in the statement of Theorem 3.1. Indeed, for this we adapt the proof of Lemma 3.3 as follows. Firstly, using Oka approximation [24, Theorem 5.4.4] (see also [7, Theorem 1.13.3]), we approximate the map $\tilde{f}_{0}\in\mathscr{A}(S,A_{*})$ in (3.2) uniformly on $S$ by $h_{0}\in\mathscr{O}(M,A_{*})$ such that $h_{0}-\tilde{f}_{0}$ vanishes to order $r_{0}+1$ on $E\cup(\Lambda\cap\mathring{S})$. Secondly, by the Mergelyan approximation theorem on admissible sets [22, Theorem 16] (see also [7, Theorem 1.12.11]), we approximate each $h_{i,j}\in\mathscr{O}(S)$ in (3.10) uniformly on $S$ by $\hat{h}_{i,j}\in\mathscr{O}(M)$ such that $\hat{h}_{i,j}-h_{i,j}$ vanishes to order $r_{0}+1$ on $E\cup(\Lambda\cap\mathring{S})$. Write $\widehat{H}=(\hat{h}_{i,j})_{i,j}$. The map $\Phi_{\widehat{H},h_{0}}\colon U\times M\to A_{*}$, defined as (3.12) but using $h_{0}$ and $\hat{h}_{i,j}$ from above, satisfies $\Phi_{\widehat{H},h_{0}}(0,\cdot)=h_{0}$, $\Phi_{\widehat{H},h_{0}}\approx\Phi_{H,\tilde{f}_{0}}$ uniformly on $U\times S$, and their difference vanishes to order $r_{0}+1$ on $U\times(E\cup(\Lambda\cap\mathring{S}))$. By the same argument as before, provided that all approximations are sufficiently close, the implicit function theorem furnishes us with a parameter $\zeta_{0}\in U$ close to the origin, such that the full map $\tilde{f}:=\Phi_{\widehat{H},h_{0}}(\zeta_{0},\cdot)/g\in\mathscr{O}_{\infty}(M|E,A_{*})$ satisfies $\mathscr{P}(\tilde{f})=\mathscr{P}(f)=0$, $\tilde{f}\approx f$ on $S$, and the difference $\tilde{f}-f$ vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$. The immersion $G$ is then obtained as in the proof of Theorem 3.1.
@@
Let $F$ be as in the statement. Set $\varepsilon_{0}:=\varepsilon/2$ and fix a nowhere vanishing holomorphic $1$-form $\theta$ on $M$. Remark 3.6 provides a neighbourhood $U\subset M$ of $S$ with $H_{1}(S,\mathbb{Z})\overset{\sim}{=}H_{1}(U,\mathbb{Z})$, and a full meromorphic $A$-immersion $F^{0}\in\mathscr{I}_{A}(U|E)$ such that $F^{0}-F$ is continuous on $S$, vanishes on $\Lambda\setminus\mathring{S}$, vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$ and $F^{0}-F||_{S}<\varepsilon_{0}$.
Choose a strongly subharmonic Morse exhaustion function $\tau\colon M\to\mathbb{R}$ with $S\subset\{\tau<0\}\subset\{\tau\leq 0\}\subset U$; see e.g. [7, Proposition 1.12.5]. We may assume that 0 is a regular value of $\tau$ and that every level set $\{\tau=c\}$, $c>0$, contains at most one critical point of $\tau$. Choose a strictly increasing sequence $0=c_{0}<c_{1}<c_{2}<\dots$ with $\lim_{j\to\infty}c_{j}=+\infty$ and such that for every $j\in\mathbb{Z}_{+}$, $c_{j}$ is a regular value of $\tau$, and the set $A_{j}=\{c_{j-1}<\tau<c_{j}\}$ ($j\in\mathbb{N}$) contains at most one critical point of $\tau$. Set $M_{j}=\{\tau\leq c_{j}\}$, $j\in\mathbb{Z}_{+}$. We shall inductively construct a sequence of full meromorphic $A$-immersions $\{F^{j}\colon M_{j}\setminus E\to\mathbb{C}^{n}\}_{j}$, $F^{j}\in\mathscr{I}_{A}(M_{j}|E)$, and a decreasing sequence of positive numbers $\{\varepsilon_{j}\}_{j}$, satisfying the following conditions for all $j\in\mathbb{N}$.
@@
    $0<\varepsilon_{j}<\varepsilon_{j-1}/2$, and if $G\colon M\setminus E\to\mathbb{C}^{n}$ is holomorphic with $G-F^{j}||_{M_{j}\setminus E}<\varepsilon_{j-1}$, then $G|_{M_{j-1}\setminus E}\colon M_{j-1}\setminus E\to\mathbb{C}^{n}$ is an immersion.
If such sequences exist, then the sequence $\{F^{j}\}_{j}$ converges uniformly on compact sets to the limit map $G:=\lim_{j\to\infty}F^{j}\colon M\setminus E\to\mathbb{C}^{n}$, which is a full meromorphic $A$-immersion of class $\mathscr{I}_{A}(M|E)$, the difference GFG-F extends continuously to $S$, vanishes on $\Lambda\setminus\mathring{S}$ and vanishes to order $r$ on $E\cup(\Lambda\cap\mathring{S})$, and $G-F||_{S}<\varepsilon$. So, $G$ satisfies the conclusion of the theorem.
@@
Case 2: The domain $A_{j}=\{c_{j-1}<\tau<c_{j}\}$ contains a critical point $p$ of $\tau$. By assumption, $p$ is the unique critical point of $\tau$ on $A_{j}$ and it is a Morse point with Morse index either 0 or $1$. If the Morse index of $\tau$ at $p$ equals $1$, the change of topology is described by attaching to $M_{j-1}$ a smoothly embedded oriented arc $\Gamma_{j}\subset A_{j}\cup bM_{j-1}$ with an initial point $p_{j}\in bM_{j-1}$ and a final point $q_{j}\in bM_{j-1}$, and such that $\Gamma_{j}$ intersects $bM_{j-1}$ transversely and is otherwise contained in $A_{j}$. The set $S_{j}:=M_{j-1}\cup\Gamma_{j}$ is admissible in $M_{j}$ and has the same topology as $M_{j}$. If the points $p_{j},q_{j}$ belong to the same connected component of $M_{j-1}$, a new nontrivial curve appears in the homology, however, if they belong to different components of $M_{j-1}$, we have that $H_{1}(M_{j-1},\mathbb{Z})\overset{\sim}{=}H_{1}(S_{j},\mathbb{Z})$. In either case, we proceed as follows. By [7, Lemma 3.5.4], we extend $f^{j-1}$ smoothly to the arc $\Gamma_{j}$ such that it still maps to $A_{*}$ and satisfies
@@
Otherwise, the Morse index of $\tau$ at $p$ equals 0. This time, a new connected component of $\{\tau\leq c\}$ appears at $p$ when $c$ passes $\tau(p)$. Let $\Delta\subset A_{j}$ denote a smoothly bounded disc around $p$. We extend $F^{j-1}$ to $S_{j}:=M_{j-1}\cup\Delta$ by defining any full holomorphic $A$-immersion on $\Delta$. Since $S_{j}$ is admissible in $M_{j}$ and has the same topology as $M_{j}$, Remark 3.6 applied to each component of $S_{j}$ furnishes a suitable map $F^{j}\in\mathscr{I}_{A}(M_{j}|E)$. This closes the induction and completes the proof. ∎
@@ 4. Desingularizing meromorphic $A$-immersions
In this section we prove the following more precise version of Corollary 1.7-(b) for removing self-intersections of directed meromorphic immersions. We shall later use it in proofs of Theorems 5.1 and 7.1 in order to ensure the embeddedness.
@@
If $M$ is a compact bordered Riemann surface, $E\subset\mathring{M}$ is a finite subset, and $A\subset\mathbb{C}^{n}$ $(n\geq 3)$ is a closed irreducible conical subvariety which is not contained in any hyperplane and is smooth away from 0, then every meromorphic $A$-immersion $F\colon M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ can be approximated uniformly on $M\setminus E$ by meromorphic $A$-embeddings $G\in\mathscr{I}_{A}(M|E)$ such that the difference map GFG-F is continuous on $M$ and vanishes to any given finite order on $E$.
@@
We emphasize that the approximation by embeddings is uniform in the non-compact surface $M\setminus E$. We establish the theorem by a recursive application of the following result.
@@
Let $M$, $E$ and $A$ be as in Theorem 4.1. Given an open neighbourhood $\Delta\subset\mathring{M}$ of $E$, every meromorphic $A$-immersion $F\colon M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ can be approximated uniformly on $M\setminus E$ by a meromorphic $A$-immersion $G\in\mathscr{I}_{A}(M|E)$ such that $G|_{M\setminus\Delta}:M\setminus\Delta\to\mathbb{C}^{n}$ is injective, the difference map GFG-F is continuous on $M$ and vanishes to any given finite order on $E$.
@@
The key condition in the lemma is that, while the self-intersections are eliminated only from a compact subset, namely, $M\setminus\Delta$, the approximation is uniform on $M\setminus E$. The proof relies on the classical Abraham transversality argument [1][7, Theorem 1.4.3] together with the method of sprays.
@@
(4.1) $\bigcup_{j\geq 1}\left(M\setminus\Delta_{j}\right)=M\setminus E.$
@@
    $F^{j}-F^{j-1}$ is continuous on $\widetilde{M}$ and $F^{j}-F^{j-1}||_{\widetilde{M}}<\varepsilon_{j-1}$.
@@
    $0<\varepsilon_{j}<\varepsilon_{j-1}/2$, and if $G\colon\widetilde{M}\setminus E\to\mathbb{C}^{n}$ is a holomorphic map with $G-F^{j}||_{\widetilde{M}\setminus E}<2\varepsilon_{j}$, then $G$ is an embedding on $M\setminus\Delta_{j}$.
@@
By (1j), (2j), (4j) and the properties of $\Delta_{j}$, there exists the limit map $G=\lim_{j\to\infty}F^{j}\colon\widetilde{M}\setminus E\to\mathbb{C}^{n}$, which is meromorphic on $\widetilde{M}$ with effective poles precisely in $E$ and satisfies that GFG-F is continuous on $\widetilde{M}$. In fact, for each $j\geq 0$ we have that $G-F^{j}$ is continuous on $\widetilde{M}$ and
@@
As a consequence, by property (4j), $G\in\mathscr{I}_{A}(M|E)$ and is injective on $M\setminus\Delta_{j}$ for all $j\geq 1$. By (4.1), it turns out that $G$ is a meromorphic $A$-embedding on $M$. Finally, condition (2j) ensure that GFG-F vanishes to the given finite order on $E\cup\Lambda$, whereas the approximation of $F$ by $G$ is guaranteed by (4.2), closing the proof. ∎
@@
The proof follows closely the arguments in that of [4, Theorem 2.4] and [7, Theorem 3.4.1] for the case $E=\varnothing$. We shall only point out the modifications which are necessary to deal with the poles and explain the first part of the lemma; the second part concerning interpolation on $\Lambda$ is established exactly as in the proof of [7, Theorem 3.4.1] and we leave it out.
@@
Let $D_{\widetilde{M}}=\{(p,p)\colon p\in\widetilde{M}\}$ be the diagonal of $\widetilde{M}\times\widetilde{M}$. Since $F$ is an immersion on $M\setminus E$ and $\widetilde{M}\subset M\setminus E$ is compact, there exists an open neighbourhood $U\subset\widetilde{M}\times\widetilde{M}$ of $D_{\widetilde{M}}$ such that $\delta F$ does not take the value $0\in\mathbb{C}^{n}$ on $\overline{U}\setminus D_{\widetilde{M}}$. Our aim is to find a meromorphic $A$-immersion $G\colon M\setminus E\to\mathbb{C}^{n}$ of class $\mathscr{I}_{A}(M|E)$ such that GFG-F is continuous and arbitrarily close to 0 uniformly on $M$, GFG-F vanishes to the given finite order on $E$, and the difference map $\delta G$ is transverse to $0\in\mathbb{C}^{n}$ on $(\widetilde{M}\times\widetilde{M})\setminus U$. By dimension reasons and the Abraham transversality argument [1], provided that $G$ is close enough to $F$ on $\overline{U}$, then $G$ results to be injective on $\widetilde{M}=M\setminus\Delta$. For this, we need to find a neighbourhood $\Omega\subset\mathbb{C}^{N}$ of $0\in\mathbb{C}^{N}$ (for some large $N\in\mathbb{N}$) and a holomorphic map $H\colon(M\setminus E)\times\Omega\to\mathbb{C}^{n}$ meeting the following.
@@
    $\partial_{t}|_{t=0}\delta H\left(p,q,t\right)\colon\mathbb{C}^{N}\longrightarrow\mathbb{C}^{n}$
@@
That being the case, the map $G=H(\cdot,t)$ satisfies the conclusion for almost every parameter $t\in\Omega$ close enough to 0.
@@
For any pair of points $\in M\setminus E, pqp\neq q$, there is a deformation family $H=H^{(p,q)}(\cdot,t)$ as above, $t\in\mathbb{C}^{n}$, such that $\partial_{t}|_{t=0}\delta H(p,q,t):\mathbb{C}^{n}\to\mathbb{C}^{n}$ is an isomorphism.
@@
(cf. [4, Eq. (6.6)]), it follows that $H_{F}(\cdot,0)=F$ and $H_{F}(\cdot,t):M\setminus E\to\mathbb{C}^{n}$ is a well-defined meromorphic $A$-immersion of class $\mathscr{I}_{A}(M|E)$ such that $H_{F}(\cdot,t)-F$ is continuous on $M$ and vanishes to the given finite order on $E$ for every $t\in\mathbb{C}^{n}$ sufficiently close to 0. Furthermore, we may ensure that the vectors
@@
defined analogously to those in the proofs of [4, Theorem 2.4] and [7, Theorem 3.4.1], satisfies $H(\cdot,0)=F$ and its difference map $\delta H$ (4.3) is submersive on $(\widetilde{M}\times\widetilde{M})\setminus U$ for all $t=(t^{1},\dots,t^{m})$ close to $0\in\mathbb{C}^{N}$ (here, $N=mn$). Moreover, $H(\cdot,t)$ is a meromorphic $A$-immersion of class $\mathscr{I}_{A}(M|E)$, $H(\cdot,t)-F$ is continuous on $M$ and vanishes to any given finite order on $E$ for every $t\in\Omega\subset\mathbb{C}^{N}$ (where $\Omega\subset\mathbb{C}^{N}$ is a neighbourhood of 0 which depends on all $\Omega_{i}$).